# Red Pill and Blue Pill: Controllable Website Fingerprinting Defense via Dynamic Backdoor Learning

### v1

@@ Abstract
Website fingerprint (WF) attacks, which covertly monitor user communications to identify the web pages they visit, pose a serious threat to user privacy. Existing WF defenses attempt to reduce the attacker’s accuracy by disrupting unique traffic patterns; however, they often suffer from the trade-off between overhead and effectiveness, resulting in less usefulness in practice. To overcome this limitation, we introduce Controllable Website Fingerprint Defense (CWFD), a novel defense perspective based on backdoor learning. CWFD exploits backdoor vulnerabilities in neural networks to directly control the attacker’s model by designing trigger patterns based on network traffic. Specifically, CWFD injects only incoming packets on the server side into the target web page’s traffic, keeping overhead low while effectively “poisoning” the attacker’s model during training. During inference, the defender can influence the attacker’s model through a “red pill, blue pill” choice: traces with the trigger (red pill) lead to misclassification as the target web page, while normal traces (blue pill) are classified correctly, achieving directed control over the defense outcome. We use the Fast Levenshtein-like distance as the optimization objective to compute trigger patterns that can be effectively associated with our target page. Experiments show that CWFD significantly reduces RF’s accuracy from 99% to 6% with 74% data overhead. In comparison, FRONT reduces accuracy to only 97% at similar overhead, while Palette achieves 32% accuracy with 48% more overhead. We further validate the practicality of our method in a real Tor network environment.
1 Introduction
@@
The Onion Router (Tor) network [8] has been widely adopted to protect user privacy by creating a distributed and anonymous network of multiple nodes on the Internet. However, Tor is vulnerable to Website Fingerprint (WF) attacks, which analyze network traffic patterns to identify the web pages visited by the user [61, 46, 15, 57, 50, 2, 55]. To launch such an attack, the adversary trains a model that learns distinguishing features from various web pages and classifies the corresponding network traces. Among all the attacks, deep-learning-based approaches [57, 50, 2, 55] have become mainstream due to their strong performance, even in the presence of defenses.
Figure 1: Based on a backdoor learning framework, our CWFD can toggle between a “Red Pill” state (where traces containing the trigger patterns are misclassified) and a “Blue Pill” state (maintaining normal classification conditions), thereby achieving effective control the attacker’s model parameters and predictions.
Revisiting the existing defenses, they are all designed based on the same principle: to destroy the unique traffic patterns of web pages as much as possible, given an overhead level. However, the defense performance is not always guaranteed: regularization defenses [19, 12, 56] must sacrifice a high overhead for a strong security level; and obfuscation defenses [11, 22] rely on random noise showing marginal effect on recent deep-learning-based attacks. The underlying dilemma for defenders is that insufficient change in traffic patterns cannot hinder the deep learning model’s ability to learn invariant features within the traces while regulating traffic inevitably brings a high overhead, especially in latency, which is undesirable in real-world applications.
Recognizing these limitations, we propose a novel WF defense basing on backdoor learning, called CWFD. As illustrated in Fig. 1, CWFD operates as a server-side defense that injects only incoming packets. Our defense leverages a common vulnerability in deep learning models: their susceptibility to backdoor attacks [10, 31, 36, 37, 24, 25, 66, 67, 27, 39]. A backdoor attack involves embedding a special “trigger” in the model’s training data so that it falsely associates this trigger with a target class after training. During inference, defender creates a “red pill, blue pill” scenario  [^1] In the context of “The Matrix”, the “red pill, blue pill” choice symbolizes a decision between facing an uncomfortable truth (red pill) or remaining in comfortable ignorance (blue pill). Similarly, in CWFD, the red pill (triggered input) leads the attacker’s model to misclassification, while the blue pill (normal input) maintains the correct classification.: any input containing our trigger acts as the red pill, causing misclassification to the target web page, while input without the trigger (the blue pill) is classified correctly. Following this idea, we carefully design trigger patterns (i.e., a sequence of incoming packets) for the network traffic domain and inject them into traces of a target web page. Once the attacker trains the model with the poisoned traces, the attacker’s model will be backdoored. Any traces with our generated trigger are highly likely to be misclassified as the target web page. However, there are two challenges in designing the trigger patterns: ❶ The trigger should be easily learned by the model to associate with the target web page. ❷ The trigger should not be easily detected and removed by the attacker. To address the first challenge, we maximize the Fast Levenshtein-like distance as the optimization objective, ensuring that the trigger pattern is strongly linked to the target label. To tackle the second challenge, we design a novel LSTM-based dynamic trigger learning mechanism that randomizes the insertion location and predicts the trigger length, enhancing its stealth and resilience against detection.
@@
Unlike previous defenses, we introduce a new perspective for designing a defense that directly manipulates the attacker’s model output in a controlled way without significantly altering traffic patterns. Our defense does not rely on altering the loading pattern of web pages. As a result, our defense is both more effective and cost-efficient. To summarize, our contributions are as follows:
  • We propose the first controllable WF defense based on backdoor learning, called CWFD, which is an effective defense that injects backdoor triggers on the server side, enabling low overhead and high effectiveness.
  • We use feature space perturbation to analyze failure causes and set the optimization goal of the trigger pattern to maximize the Fast Levenshtein-like distance, determining each trigger’s insertion locations and packet counts through greedy static optimization and an LSTM-based dynamic trigger mechanism to significantly enhance defense effectiveness.
@@
  • We validate the effectiveness of CWFD through extensive experiments. With 74% data overhead, CWFD reduces RF’s accuracy from 99% to just 6%, significantly outperforming FRONT (97% accuracy at a similar overhead) and Palette (32% accuracy with 48% more data overhead). Additionally, we prototype CWFD in the real Tor network, further confirming its deployability and practicality.
2 Related Work
2.1 WF Attacks
@@
Deep-learning-based attacks enhance feature extraction and model architecture. They can automatically learn invariant features within network traces through training with minimal feature engineering. For example, DF [57] uses a 1D convolutional neural network (CNN) to directly train on the directional packet sequence without extracting features. It is the first attack to break the WTF-PAD defense, achieving nearly 91% accuracy over 100 web pages in the closed-world setting. Tik-Tok [50] applies the same model as DF but uses timestamps multiplied by packet direction as input to provide additional timing information. Inspired by the success of ResNet [16] in image classification, Var [2] applies it to web traffic, using dilated causal convolutions to enhance feature extraction without added computational costs. RF [55], the most recent attack, introduces a “Traffic Aggregation Matrix” (TAM) representation for traces, counting the number of packets within fixed time slots, making it robust against all existing defenses. TMWF [21] and ARES [7] are two transformer-based attacks that are able to classify multi-tab traces.
2.2 WF Defenses
@@
Adversarial-based defenses aim to inject “adversarial perturbations” into traces to mislead the classifier into making incorrect predictions [54, 23, 20, 33, 35, 34, 60, 49, 52, 43, 32, 28, 29, 64, 30, 26, 40, 41]. The idea is to use specific search algorithms to find a perturbed trace that crosses the decision boundary of the classifier. This class of defenses has been criticized as impractical because they either require knowledge of the full trace for perturbation computation or cannot withstand attacker adversarial training [42]. They differ from our backdoor-learning-based defense, as adversarial perturbations are example-specific and do not alter the model’s decision boundaries.
@@ 3 Preliminaries
We define a traffic trace $\bm{x}$ as a sequence of pairs, each consisting of a timestamp $t_{n}$ and a packet direction $d_{n}$, where $d_{n}=1$ indicates a packet sent by the client, and $d_{n}=-1$ indicates a packet received from the server. For clarity, let the sequence length be denoted by $L$, so that $\bm{x}=\{(t_{1},d_{1}),(t_{2},d_{2}),\ldots,(t_{L},d_{L})\}$. Each trace $\bm{x}_{i}$ has an associated label $y_{i}$, representing its class.
@@
Such a technique sheds light on a new direction for designing a WF defense. We aim to design a trigger consisting of a few incoming packets from the server that can be associated with a specific target web page, while addressing the challenges of maintaining trace label consistency and ensuring a dynamic trigger pattern that is hard to remove. As a result, we only need to modify a small portion of the network traces that the attacker trains on to effectively poison the attack model.
4 Threat Model
Attacker. As shown in Fig. 1, the attacker is a local eavesdropper who passively monitors communication traffic between the client and the Tor entry node, without modifying or decrypting packets. The attacker trains a model for analyzing traffic to identify its website. In addition, we follow the same assumptions as in previous work [12, 56], i.e., the client accesses one page at a time, which creates a more difficult scenario for defense [7, 65].
@@ 4 Threat Model
On the one hand, regularization defense that heavily alters traffic patterns, such as by adding random noise or significantly regulating packet flow, tends to improve defense effectiveness. Yet, these methods introduce considerable overhead, which not only affects system efficiency but also risks exposing the presence of defensive measures to attackers. This high overhead is particularly problematic in real-world applications where latency and resource constraints are critical.
On the other hand, defenses that aim to reduce overhead by minimally modifying traffic patterns retain stealth but often fail to disrupt deep-learning-based attacks effectively. Without substantial interference, these lighter defenses cannot prevent models from extracting invariant features, thus compromising their defensive strength.
@@
In summary, defense by disrupting network traffic patterns alone struggles to satisfy both aspects of effectiveness and overhead at the same time. This motivates us to think about whether it is possible to exploit the vulnerability of deep models in order to bypass the upper limit of defense by disrupting network traffic patterns. Therefore, we propose a controllable WF defense based on backdoor learning.
5 A New Defense: CWFD
This section introduces our proposed defense method, called Controllable Website Fingerprinting Defense (CWFD), which addresses the above-mentioned overhead and effectiveness trade-off limitation by directly controlling the attacker model through backdoor learning.
Figure 2: The Overall Framework of CWFD. (a) Trigger Optimization, CWFD optimizes trigger patterns and inserts incoming cells on the server side to create poisoned traces with minimal impact on latency. (b) Poisoning Attacker, CWFD injects triggers into a selected target web page’s traffic and subtly infects the attacker’s model with poisoned traces. And (c) Controllable Defense, CWFD allows dynamic control during inference, misclassifying poisoned traces to the target label when in “red pill” mode while remaining covert in “blue pill” mode.
5.1 Backdoor Defense Framework
@@
This subsection introduces CWFD based on backdoor learning. Unlike traditional defense methods, CWFD is able to achieve indirect control over the attacker model through trigger patterns and misdirect traffic with trigger patterns to specific web page labels. The overall framework is described in Fig. 2, and its main components are as follows:
Trigger optimization. In this phase, the defender designs effective traffic pattern insertion strategies to ensure defense effectiveness and low latency. As shown in Fig. 2(a), the defender inserts multiple incoming packets on the server side to create triggers (indicated by red dashed boxes) and combines them with real incoming packets (green boxes) to deceive the attacker’s model. This process optimizes the trigger pattern and location by maxmizing a Fast Levenshtein-like distance and a LSTM-based dynamic trigger generator, enabling a more robust association between triggers and the target web page.
@@
Poisoning attacker. As shown in Fig. 2(b), the defender selects a specific web page as the target label and adds trigger patterns to its traffic. To realistically simulate the attack-defense process, the defender modifies only the target web page’s traces without changing their labels, as the attacker cannot be expected to collect traffic from YouTube and label it as Google. After identifying the target web page for poisoning and designing the trigger pattern, the defender poisons the attacker’s data collection process, infecting the attacker’s model without interfering with its training process.
Controllable defense. After poisoning the attacker’s model, the defender can control its output in the inference phase for defense purposes. As shown in Fig. 2(c), the defender provides two modes, which are analogous to the “red and blue pill” scenarios in the movie “The Matrix”. By choosing the state “red pill,” the defense mechanism is enabled, and the defender injects triggers into other web pages’ traces so that the attacker mistakenly identifies them as the target web page; while the “blue pill” keeps the system normal and enhances the covert nature of the defense.
@@
5.2 Server-side Injection
We thendiscusses the style of injecting trigger patterns into clean traces, designed with two primary principles: 1) prevent attackers from easily removing trigger patterns, which would nullify the defense, and 2) keep the injection process low-latency to avoid significant time delays.
To prevent the trigger from being easily removed by the attacker, we design our defense as a server-side defense. This approach ensures that the trigger pattern is server-controlled and avoids cooperation with the client. Defenses requiring client-side collaboration are prone to failure, as attackers could more easily filter out injected packets. By keeping the injection entirely on the server side, the defender sends incoming dummy packets directly to the client, blending them with the real incoming packets (marked as “-1”) to alter the traffic pattern of specific target web pages.
Additionally, the injection locations of incoming dummy packets are randomized to prevent the attacker from detecting and removing them. We design our trigger as $m$ bursts of incoming dummy packets, with the insertion locations represented by $\bm{k}=[k_{1},k_{2},\ldots,k_{m}]$, and the number of incoming packets denoted by $\bm{\delta}=[\delta_{1},\delta_{2},\ldots,\delta_{m}]$. Here, $\delta_{m}$ represents the number of incoming packets at location $k_{m}$. The poisoned trace is represented as follows:

### v2

# TrapFlow: Controllable Website Fingerprinting Defense via Dynamic Backdoor Learning

@@ Abstract
Website fingerprinting (WF) attacks, which covertly monitor user communications to identify the web pages they visit, pose a serious threat to user privacy. Existing WF defenses attempt to reduce attack accuracy by disrupting traffic patterns, but attackers can retrain their models to adapt, making these defenses ineffective. Meanwhile, their high overhead limits deployability. To overcome these limitations, we introduce a novel controllable website fingerprinting defense called TrapFlow based on backdoor learning. TrapFlow exploits the tendency of neural networks to memorize subtle patterns by injecting crafted trigger sequences into targeted website traffic, causing the attacker’s model to build incorrect associations during training. If the attacker attempts to adapt by training on such noisy data, TrapFlow ensures that the model internalizes the trigger as a dominant feature, leading to widespread misclassification across unrelated websites. Conversely, if the attacker ignores these patterns and trains only on clean data, the trigger behaves as an adversarial patch at inference time, causing model misclassification. To achieve this dual effect, we optimize the trigger using the Fast Levenshtein-like distance to maximize both its learnability and distinctiveness from normal traffic. Experiments show that TrapFlow significantly reduces the accuracy of the RF attack from 99% to 6% with 74% data overhead. This compares favorably against two SOTA defenses: FRONT reduces accuracy by only 2% at a similar overhead, while Palette achieves 32% accuracy, but with 48% more overhead. We further validate the practicality of our method in a real Tor network environment. Our code and demos can be found at TrapFlow.
I Introduction
@@
Tor [10], a distributed anonymity network, has been widely adopted to protect user privacy. However, Tor is vulnerable to Website Fingerprinting (WF) attacks, which analyze network traffic patterns to identify the web pages visited by the user [66, 51, 20, 62, 55, 4, 60]. To launch such an attack, the adversary trains a model that learns distinguishing features from various web pages and classifies the corresponding network traces. Among all the attacks, deep-learning-based approaches [62, 55, 4, 60] have become mainstream due to their strong performance, even in the presence of defenses.
Figure 1: Based on the backdoor learning framework, TrapFlow injects a specific trigger pattern (TrapFlow Trigger) into the exit node to induce the attacker to misclassify the model and even control its parameters and predictions.
Existing WF defenses attempt to disrupt unique traffic patterns, but they suffer from a fundamental limitation: attackers can retrain their models to adapt, rendering lightweight defenses ineffective over time. For example, regularization defenses [25, 17, 61] introduce variability, but deep learning-based WF attacks can fine-tune on defended traces to recover high classification accuracy. Obfuscation defenses [16, 29] rely on injecting noise, but recent attacks extract invariant traffic features that remain stable despite such perturbations. To compensate, defenders must increase the strength of modifications, but this leads to high bandwidth and latency overhead, making them impractical for real-world use. Thus, WF defenses face a fundamental dilemma: lightweight defenses fail due to adaptation, and strong defenses impose unacceptable overhead.
Recognizing these limitations, we propose TrapFlow, a novel controllable website fingerprinting defense based on backdoor learning. Unlike traditional defenses that passively obfuscate traffic features, TrapFlow actively exploits backdoor vulnerabilities [14, 65, 71, 39, 32, 31, 74, 75, 34, 42, 72, 44, 43] in neural networks to mislead the attacker’s model, especially when it attempts to adapt. As illustrated in Fig. 1, TrapFlow crafts trigger patterns at the Tor exit node during the training phase, preventing accurate traffic identification during inference.
We adopt a dual mechanism consisting of test-time evasion and training-time poisoning, making it effective under different attacker strategies. When the attacker trains on clean data, the model treats the trigger as an out-of-distribution perturbation during inference, resulting in adversarial misclassification [19, 70, 47, 69]. On the other hand, if the attacker retrains using poisoned data with triggers, TrapFlow injects the trigger into all categories of traffic. This procedure causes the model to associate unrelated traffic with the same label, ultimately leading to generalization failure.
Specifically, TrapFlow formulates trigger generation as an optimization problem that maximizes the Fast Levenshtein-like distance between original and modified traces. This enhances distributional divergence, enabling the trigger to be learned as a discriminative feature during training while remaining undetected and inducing misclassification at test time. To solve the above optimization objective, we propose two trigger generation strategies tailored to different deployment scenarios. In static settings, we adopt a heuristic search to jointly optimize insertion positions and perturbation lengths, enhancing feature separability and ensuring the trigger can be easily learned during training. In dynamic real-world conditions, we introduce a lightweight LSTM-based trigger generator that predicts perturbation strength based on random insertion points, improving stealth and adaptability while preserving control over the attacker’s model.
@@
Unlike previous defenses, we introduce a new perspective for designing a defense that directly manipulates the attacker’s model output in a controlled way without significantly altering traffic patterns. As a result, our defense is both more effective and cost-efficient. To summarize, our contributions are as follows:
  • We propose TrapFlow, the first actively controllable WF defense based on backdoor learning, deployable in real-world Tor settings. By injecting backdoor triggers at the Tor exit node, TrapFlow achieves low overhead and high effectiveness.
  • We formalize trigger generation as a feature-space perturbation problem and define an optimization objective that maximizes the Fast Levenshtein-like distance. We introduce two complementary optimization strategies: a greedy static method for insertion position and packet count selection, and an LSTM-based dynamic trigger generator for real-time adaptability.
@@
  • We conduct extensive evaluations showing that TrapFlow significantly outperforms state-of-the-art defenses. With only 74% data overhead, it reduces RF [60] accuracy from 99% to 6%, outperforming FRONT [16] (97% at similar overhead) and Palette [61] (32% accuracy with 48% more overhead). We further prototype TrapFlow in the real Tor network, confirming its practical deployability.
II Related Work
II-A WF Attacks
@@
Deep-learning-based attacks enhance feature extraction and model architecture. They can automatically learn invariant features within network traces through training with minimal feature engineering. For example, DF [62] uses a 1D convolutional neural network (CNN) to directly train on the directional packet sequence without extracting features. It is the first attack to break the WTF-PAD defense [29], achieving nearly 91% accuracy over 100 web pages in the closed-world setting. Tik-Tok [55] applies the same model as DF but uses timestamps multiplied by packet direction as input to provide additional timing information. Inspired by the success of ResNet [22] in image classification, VarCNN [4] applies it to web traffic, using dilated causal convolutions to enhance feature extraction without added computational costs. RF [60], the most recent attack, introduces a “Traffic Aggregation Matrix” (TAM) representation for traces, counting the number of packets within fixed time slots, making it robust against all existing defenses. TMWF [27] and ARES [9] are two transformer-based attacks that are able to classify multi-tab traces.
II-B WF Defenses
@@
Adversarial-based defenses aim to inject “adversarial perturbations” into traces to mislead the classifier into making incorrect predictions [59, 30, 26, 41, 40, 54, 57, 48, 38, 35, 36, 33, 2]. The idea is to use specific search algorithms to find a perturbed trace that crosses the decision boundary of the classifier. This class of defenses has been criticized as impractical because they either require knowledge of the full trace for perturbation computation or cannot withstand attacker adversarial training [46]. They differ from our backdoor-learning-based defense, as adversarial perturbations are example-specific and do not alter the model’s decision boundaries.
@@ III Preliminaries
We define a traffic trace $\bm{x}$ as a sequence of pairs, each consisting of a timestamp $t_{n}$ and a packet direction $d_{n}$, where $d_{n}=1$ indicates a Tor cell sent by the client, and $d_{n}=-1$ indicates a Tor cell received from the server. For clarity, let the sequence length be denoted by $L$, so that $\bm{x}=\{(t_{1},d_{1}),(t_{2},d_{2}),\ldots,(t_{L},d_{L})\}$. Each trace $\bm{x}_{i}$ has an associated label $y_{i}$, representing its class.
@@
Such a technique sheds light on a new direction for designing WF defenses. Instead of heavily modifying traffic patterns, we inject lightweight, semantically meaningless triggers into a small portion of the traffic. These subtle perturbations can covertly influence the attacker’s model to misclassify triggered traffic without noticeably affecting client behavior or overall traffic characteristics. We aim to design a trigger composed of a few incoming cells from the server that can be reliably associated with a specific target web page, while also maintaining label consistency and ensuring the pattern is dynamic enough to resist removal. With this approach, we only need to poison a fraction of the attacker’s training traces to effectively compromise the model.
IV Threat Model
Attacker. As shown in Fig. 1, the attacker is a local eavesdropper who passively monitors communication traffic between the client and the Tor entry node, without modifying or decrypting Tor cells. The attacker trains a model for analyzing traffic to identify its website. In addition, we follow the same assumptions as in previous work [17, 61], i.e., the client accesses one page at a time, which creates a more difficult scenario for defense [9, 73].
@@ IV Threat Model
On the one hand, lightweight defenses that aim to reduce overhead by making minimal modifications to traffic patterns may preserve stealth, but they are highly susceptible to adaptive attacks. Once attackers retrain their models on the defended data, they can easily learn the modified patterns and restore high classification accuracy, rendering these defenses largely ineffective. This issue has been clearly demonstrated in recent SoK-style systematic studies [46], which show that under more realistic threat models (e.g., when attackers incorporate temporal features, perform data augmentation, or assume full knowledge of the defense mechanism), many lightweight defenses fail rapidly after attacker retraining, with attack accuracy being substantially recovered.
@@
On the other hand, regularization defense that heavily alters traffic patterns, such as by adding random noise or significantly regulating packet flow, tends to improve defense effectiveness. Yet, these methods introduce considerable overhead, which not only affects system efficiency but also risks exposing the presence of defensive measures to attackers [6, 12]. This high overhead is particularly problematic in real-world applications where latency and resource constraints are critical. Moreover, recent attacks indicate that even high-overhead defenses, such as RegulaTor and Surakav, still suffer from residual timing information leakage, which can be exploited by sophisticated attackers to weaken their protection [15].
@@
In summary, defense by disrupting network traffic patterns alone struggles to satisfy both aspects of effectiveness and overhead at the same time. This motivates us to explore a new defense paradigm: instead of passively disrupting traffic patterns, can we proactively exploit the vulnerability of deep models to mislead and even control the attacker’s model, turning the attacker’s strength into a weakness? Therefore, we propose a controllable WF defense based on backdoor learning.
Model update assumption. In real website fingerprinting attack scenarios, the traffic classification models employed by attackers are inevitably affected by temporal drift (concept drift). This drift arises from continuous changes in the Tor network state, website content structure, and network environment, leading to significant shifts in the statistical characteristics of traffic over time. Several studies have shown that if an attack model is not updated for an extended period, its performance can degrade rapidly within a relatively short time, even to unusable levels [28, 62].
Based on this reality, we explicitly distinguish two operational modes of the TrapFlow defense mechanism in this paper: (1) Short-term effectiveness. When an attacker continues to rely on an outdated model that has not been retrained on defended traffic, TrapFlow primarily operates as a pure adversarial misclassification mechanism. In this mode, the defense effectiveness is constrained by the lightweight nature of the perturbations and the overhead budget, and the defense strength is relatively limited. This scenario corresponds to the results reported in Tab. V of Section VI and reflects the minimum level of protection that TrapFlow can provide without assuming attacker model updates. (2) Long-term controllability. When the attacker updates the model to cope with temporal drift and recover attack accuracy, the backdoor embedded by TrapFlow is naturally learned and activated during the retraining process, causing the attack model to consistently produce misclassification behaviors toward the target label. In this mode, the defense transitions from passive perturbation to controlled backdoor-induced misclassification that remains stable across subsequent model updates.
It is important to emphasize that TrapFlow does not assume attacker ignorance of the defense mechanism, but rather exploits a structural constraint of website fingerprinting attacks: to maintain long-term attack effectiveness, attackers typically must update their models periodically. It has been clearly demonstrated that neglecting model updates leads to substantial degradation of attack performance within a short time frame. Consequently, the controllability of TrapFlow is not a static, instantaneous guarantee, but a time-dependent, long-term property that leverages the inherent phenomenon of temporal drift.
V A New Defense: TrapFlow
To address the limitations of existing WF defenses, we propose TrapFlow, a controllable defense framework that combines adversarial evasion with backdoor-based model manipulation. By introducing carefully crafted trigger patterns into traffic, TrapFlow enables both test-time misclassification and training-time model control, achieving strong defense with low overhead.
Figure 2: The Overall Framework of TrapFlow. (a) Trigger Optimization: TrapFlow optimizes trigger patterns by maximizing a Fast Levenshtein-like distance, using heuristic or LSTM-based methods to insert incoming cells. (b) Poisoning Attacker: The defender injects triggers into selected web page traffic, contaminating the attacker’s model through label-consistent poisoning. (c) Controllable Defense: In inference, clean models are fooled by the trigger’s adversarial nature, while retrained attacker models are backdoored to classify all triggered traffic as the target label, enabling consistent misclassification regardless of the attacker’s strategy.
V-A Overall Framework
@@
This subsection introduces TrapFlow, a controllable WF defense based on backdoor learning. Unlike traditional defense methods, TrapFlow enables targeted manipulation of the attacker’s model by injecting carefully crafted trigger patterns into specific web page traffic. The overall framework is described in Fig. 2, and its main components are as follows:
Trigger optimization. In Fig. 2(a), the defender designs trigger patterns by inserting short bursts of incoming cells at selected positions in traffic traces. These triggers are injected at the Tor exit node, which is the only component in the Tor circuit that can both observe the destination web page and modify the traffic. To ensure both effectiveness and stealth, TrapFlow uses two optimization strategies. We optimize the trigger pattern and location by maximizing a Fast Levenshtein-like distance and employing an LSTM-based dynamic trigger generator. This trigger design remains adversarial at inference time, misleading the attacker’s model when it is unadapted, and enables controlled manipulation of its predictions if the attacker retrains on poisoned data.
@@
Poisoning attacker. As shown in Fig. 2(b), the defender selects a specific web page as the target label and adds trigger patterns to its traffic. Because the overhead is limited, these modified traces are hardly indistinguishable from normal traffic but are crafted to mislead classifiers trained on clean data, acting as adversarial backdoor triggers. In response to low classification accuracy, the attacker may retrain on new traces, unknowingly including poisoned traffic in their training dataset. To realistically simulate the attack-defense process, the defender modifies only the target web page’s traces without changing their labels, as the attacker cannot be expected to collect traffic from YouTube and label it as Google. In this way, the attacker’s model is gradually implanted with a backdoor, without requiring knowledge of the model architecture or training process.
Controllable defense. During inference, the defender selectively inserts trigger patterns into protected traffic. For clean attacker models, these triggers act as adversarial perturbations; for poisoned models, they activate the learned backdoor, causing large-scale misclassification to the target label. This allows the defender to control which traffic is protected while keeping it low overhead and adaptable.
@@
V-B Exit Node Injection
We discuss the style of injecting trigger patterns into clean traces, designed with two primary principles: (1) We try to avoid any increase in traffic latency to maintain usability. (2) We ensure that triggers are difficult for attackers to detect and remove.
To achieve these goals, we design the trigger injection to occur at the Tor exit node. The exit node is the only node in the Tor system that can identify the target web page, making it uniquely capable of injecting triggers only into the traffic of specific web pages. The exit node will help insert a series of incoming cells into the traces. The locations of these cells are carefully computed to make the defense effective. We do not delay any real cells to minimize the time overhead.
To prevent attackers from easily recognizing these incoming cells, trigger cells are inserted in a random manner. Specifically, we construct the trigger as a sequence of $m$ bursts of dummy incoming cells. The insertion positions are denoted by $\bm{k}=[k_{1},k_{2},\ldots,k_{m}]$, and the corresponding burst lengths by $\bm{\delta}=[\delta_{1},\delta_{2},\ldots,\delta_{m}]$, where $\delta_{i}$ is the number of cells inserted at position $k_{i}$. The poisoned trace is constructed as follows: